# Order-2 Delaunay Triangulations Optimize Angles

### v3

@@ 1. Introduction
This paper is motivated by the desire to generalize optimal properties from order-$1$ to higher-order Delaunay triangulations. The classic (order-$1$) Delaunay triangulation (also called Delaunay mosaic) of a finite point set was introduced in 1934 by Boris Delaunay (also Delone). It is the edge-to-edge tiling whose polygons satisfy the empty circle criterion [5]: each polygon is inscribed in a circle and all other points lie strictly outside this circle. In the henceforth considered generic case, all tiles are triangles. The criterion implies that for an edge shared by two triangles, the sum of the two angles opposite to the edge is less than $\pi$. If a triangulation satisfies this criterion for every edge shared by two triangles, then we say the triangulation has the local angle property. Recognizing the potential of this type of triangulation for applications, Lawson in 1977 turned the empty circle criterion into an iterative algorithm that converts any triangulation of a given set of $n$ points in ${{\mathbb{R}}}^{2}$ into the Delaunay triangulation using at most $O(n^{2})$ edge-flips [13]. The correctness of this algorithm implies that the Delaunay triangulation is the only triangulation of the given set that has the local angle property. Using Lawson’s algorithm as a proof technique, Sibson proved in 1978 that among all triangulations of a finite generic point set in ${{\mathbb{R}}}^{2}$, the Delaunay triangulation lexicographically maximizes the vector whose components are the angles inside the triangles sorted in non-decreasing order [21]. We call this the sorted angle vector of the triangulation.
@@
With the exception of Eppstein’s result—which is specific to the farthest-point Delaunay triangulation—there is a paucity of optimality properties known for higher-order Delaunay trinagulations, which we end with three inter-related contributions:
@@
This paper is organized as follows. Section 2 provides information on the main background, including level-$k$ hypertriangulations (maximal, complete, and otherwise) and the aging function. Section 3 introduces our extension of the local angle property to order $k$ and, in Theorem 3.3, shows that the order-$k$ Delaunay triangulation has this property. Section 4 proves Property II in Theorem 4.5 and, in Theorem 4.4, gives a new short proof of Sibson’s theorem on angle vector optimality for the order-1 Delaunay triangulation [21]. It also discusses possible extensions to the class of maximal level-$2$ hypertriangulations and to levels beyond $2$. Section 5 proves Property III in Theorem 5.4, which it extends it to order-$3$ for points in convex position in Theorem 5.5. Finally, Section 6 concludes the paper with discussions of open questions and conjectures related to the geometry and combinatorics of Delaunay and more general hypertriangulations.
@@ 2.1. Triangulations and Hypertriangulations
Observe that every triangulation of $A$ is a level-$1$ hypertriangulation of $A$, and vice versa, but for $k>1$, only a subset of the triangulations of $A^{(k)}$ are level-$k$ hypertriangulations of $A$. Note also that it is possible that a point can be written as the average of more than one subset of $k$ points in $A$: for example, the center of a square is the $2$-fold average of two pairs of diagonally opposite vertices. If a level-$k$ hypertriangulation uses such a point as a vertex, then it can use only one of the possible labels.
An alternative approach to these concepts is via induced subdivisions; see [23, Chapter 9] for details, including the definitions of induced subdivisions and tight subdivisions. According to this approach, a triangulation of $A=\{a_{1},a_{2},\ldots,a_{n}\}$ is a tight subdivision of ${{\rm conv\,}{A}}$ induced by the projection $\pi\colon\Delta_{n}\to{{\mathbb{R}}}^{2}$, in which $\Delta_{n}={{\rm conv\,}{\{e_{1},e_{2},\ldots,e_{n}\}}}\subseteq{{\mathbb{R}}}^{n}$ is the standard $(n-1)$-simplex, and $\pi(e_{i})=a_{i}$, for $i=1,2,\ldots,n$. To generalize, Olarte and Santos [16] use the level-$k$ hypersimplex, $\Delta_{n}^{(k)}$, which is the convex hull of the $k$-fold averages of the $e_{i}$ in ${{\mathbb{R}}}^{n}$, and define a level-$k$ hypertriangulation of $A$ as a tight subdivision of $A^{(k)}$ induced by the same projection $\pi$ restricted to $\Delta_{n}^{(k)}$. In this setting, the constraint to use only one label for each vertex is implicit.
@@ 2.2. The Aging Function
The aging function increases the level of the triangle by one, hence the name. Correspondingly, the inverse aging function maps a black triangle to a white triangle one level lower.
To extend this definition to hypertriangulations, we say a level-$k$ hypertriangulation, $P_{k}$, ages to a level-$(k+1)$ hypertriangulation, $P_{k+1}$, denoted $P_{k+1}=F(P_{k})$. if the aging function defines a bijection between the white triangles in $P_{k}$ and the black triangles in $P_{k+1}$. Note however that the aging of $P_{k}$ is not unique as it says nothing about the white triangles of $P_{k+1}$. This notion is useful to obtain structural results for the family of all level-$k$ hypertriangulations. For example, [7] has shown that every level-$2$ hypertriangulation is an aging of a level-$1$ hypertriangulation. For the special case in which the points are in convex position, [10] has extended this result to all levels, $k$. However, for points in possibly non-convex position, there are obstacles to applying the aging function. An example of a level-$2$ hypertriangulation, $P_{2}$, for which $F(P_{2})$ does not exist is given in [7, 16].
For later reference, we compile several results about the relation between level-$1$ and level-$2$ hypertriangulations obtained in [7]. Given a vertex, $x$, in a triangulation, $P$, we define the star of $x$ as the union of triangles that share $x$, denoted ${{\rm st}({P,x})}$, and shrinking the star by a factor two toward $x$, we get ${[{{{\rm st}({P,x})},x}]}=\frac{1}{2}({{\rm st}({P,x})}+x)$, which is the set of midpoints between $x$ and any point $y\in{{\rm st}({P,x})}$. Observe that the shrunken star is contained in ${{\rm conv\,}{A^{(2)}}}$ iff $x$ is an interior vertex of $P$. Indeed, $x$ necessarily belongs to the shrunken star, but if $x$ is a convex hull vertex, then $x$ lies outside ${{\rm conv\,}{A^{(2)}}}$.
@@
Let $A\subseteq{{\mathbb{R}}}^{2}$ be finite and generic, and recall that every level-$1$ hypertriangulation is just a triangulation.
@@ 2.3. Maximal and Complete Hypertriangulations
The notion of maximality extends to level-$k$ hypertriangulations, while completeness does not since there are counterexamples to the existence of the aging function from level $2$ to level $3$; see Figure 8 in [7], which is based on Example 5.1 in [16].
@@
To prove the first claim, let $n={{\#}{A}}$, $h={{\#}{H}}$, and consider a level-$2$ hypertriangulation, $P_{2}$, aged from a possibly partial triangulation, $P$, with $m+h\leq n$ vertices. Note that $P$ has $22m+h-$ 2 triangles, so $P_{2}$ has the same number of black triangles.
@@ 3.1. Higher Order Delaunay Triangulations
Figure 1. The (blue) order-$2$ Delaunay triangulation drawn on top of the (black) order-$2$ Voronoi tessellation. Not all parts of the order-$2$ Voronoi tessellation are visible in the rectangular window.
@@ 3.2. Angles of Black and White Triangles
For $k=1$, there are no black triangles, so (bb) and (bw) are void. Delaunay [5] proved that the local angle property characterizes the (closest-point) Delaunay triangulation among all (complete) triangulations of a finite point set, and this was used by Lawson [13] to construct the triangulation by repeated edge flipping. Symmetrically, for $k={{\#}{A}}-1$, there are no white triangles, so (ww) and (bw) are void. Eppstein [8] proved the local angle property for the (farthest-point) Delaunay triangulation, and the convergence of the flip-algorithm implies that it is the only (not necessarily complete) triangulation of the points that has this property. The goal of this section is to extend these result to level-$k$ hypertriangulations.
@@ 3.3. All Delaunay Triangulations Have the Local Angle Property
We conjecture that the order-$k$ Delaunay triangulation is the only level-$k$ hypertriangulation with maximally many triangles that has the local angle property. For later reference, we refer to this as the Local Angle Conjecture for hypertriangulations.
@@ 3.4. Constrained Delaunay Triangulations
Given a bounded polygonal region, $R$, it is always possible to find a triangulation, $P$, of its vertices (the endpoints of its edges) that contains all edges of the region. Hence, every triangle of $P$ lies either completely inside or completely outside the region. The restriction of $P$ to $R$ consists of the triangles inside $R$, and we call this restriction a triangulation of $R$. For some choices of $P$, the restriction to $R$ looks locally like the Delaunay triangulation, namely when every edge that passes through the interior of $R$ satisfies (ww). It is not difficult to see that such choices of triangulations exist and that their restriction to $R$ is generically unique: run Lawson’s algorithm on an initial triangulation of $R$, flipping an interior edge whenever the sum of the two opposite angles exceeds $\pi$. This is the constrained Delaunay triangulation of $R$, as introduced in 1989 by Paul Chew [3], but see also [12]. A triangle $uvw$ belongs to this specific triangulation iff it is contained in $R$ and its circumcircle does not enclose any vertex that is visible from points inside the triangle. We state a weaker necessary condition for later reference.
@@
Let $R$ be a bounded polygonal region in ${{\mathbb{R}}}^{2}$, assume its vertex set is generic, and let $u,v,w$ be vertices of $R$. If the triangle $uvw$ is contained in $R$, and its circumcircle does not enclose any vertex of $R$, then $uvw$ is a triangle in the constrained Delaunay triangulation of $R$. Similarly, if the edge $uv$ is contained in $R$ but is not an edge of $R$, and it has a circumcircle that does not enclose any vertex of $R$, then $uv$ is an edge of the constrained Delaunay triangulation of $R$.
We use constrained Delaunay triangulations to decompose white regions in aged hypertriangulations. To explain, let $P$ be a complete triangulation of a finite and generic set, $A\subseteq{{\mathbb{R}}}^{2}$, let $x\in A$ be a vertex of this triangulation, call ${{\rm wh}({P,x})}={{\rm st}({P,x})}\cap{{\rm conv\,}{(A\setminus\{x\})}}$ the white region of $x$ in $P$, and let $P(x)$ be a triangulation of ${{\rm wh}({P,x})}$. Note that ${{\rm wh}({P,x})}={{\rm st}({P,x})}$ if $x$ is an interior vertex, and ${{\rm wh}({P,x})}\subsetneq{{\rm st}({P,x})}$ if $x$ is a convex hull vertex. In the special case in which $P$ is the order-$1$ Delaunay triangulation and $P(x)$ is the constrained Delaunay triangulation of ${{\rm wh}({P,x})}$ for each $x\in A$, these sets contains all white triangles in the order-$2$ Delaunay triangulation, albeit the latter are only half the size.
More generally, we use the constrained Delaunay triangulations of the white regions to disambiguate the aging function. This is done extensively in the proofs of our main results in Sections 4 and 5.
@@ 4.2. Elementary Lemmas
Assume without loss of generality that $L$ is horizontal, and let $A^{\prime}\subseteq A$ contain all points strictly above $L$. The boundary of ${{\rm conv\,}{A}}$ is a closed convex curve, $\gamma$, and we write $\gamma^{\prime}\subseteq\gamma$ for the vertices and edges strictly above $L$. Every point $a\in A^{\prime}$ is either a vertex of $\gamma^{\prime}$, or there is an edge $ab$ in $P$, with $b$ above $L$ and further from $L$ than $a$. Hence, $ab\in Q$. We can therefore trace a path from $a$ that eventually reaches a vertex of $\gamma^{\prime}$ in $Q$, which implies that the part of $Q$ strictly above $L$ is either empty or connected. Symmetrically, the part of $Q$ strictly below $L$ is either empty or connected, which implies the claim. ∎
@@
    If $v=b$ is the only shared vertex between $abc$ and $uvw$, then $uw$ cannot cross $ab$ and $bc$.

### v4

@@ 1. Introduction
This paper is motivated by the desire to generalize optimal properties from order-$1$ to higher-order Delaunay triangulations. The classic (order-$1$) Delaunay triangulation (also called Delaunay mosaic) of a finite point set was introduced in 1934 by Boris Delaunay (also Delone). It is the edge-to-edge tiling whose polygons satisfy the empty circle criterion [6]: each polygon is inscribed in a circle and all other points lie strictly outside this circle. In the henceforth considered generic case, all polygons are triangles. The criterion implies that for an edge shared by two triangles, the sum of the two angles opposite to the edge is less than $\pi$. If a triangulation satisfies this criterion for every edge shared by two triangles, then we say the triangulation has the local angle property. Recognizing the potential of this type of triangulation for applications, Lawson in 1977 turned the empty circle criterion into an iterative algorithm that converts any triangulation of a given set of $n$ points in ${{\mathbb{R}}}^{2}$ into the Delaunay triangulation using at most $O(n^{2})$ edge-flips [14]. The correctness of this algorithm implies that the Delaunay triangulation is the only triangulation of the given set that has the local angle property. Using Lawson’s algorithm as a proof technique, Sibson proved in 1978 that among all triangulations of a finite generic point set in ${{\mathbb{R}}}^{2}$, the Delaunay triangulation lexicographically maximizes the vector whose components are the angles inside the triangles sorted in non-decreasing order [22]. We call this the sorted angle vector of the triangulation.
@@
With the exception of Eppstein’s result—which is specific to the farthest-point Delaunay triangulation—there is a paucity of optimality properties known for higher-order Delaunay triangulations, which we end with three inter-related contributions:
@@
This paper is organized as follows. Section 2 provides information on the main background, including level-$k$ hypertriangulations (maximal, complete, and otherwise) and the aging function. Section 3 introduces our extension of the local angle property to order $k$ and, in Theorem 3.3, shows that the order-$k$ Delaunay triangulation has this property. Section 4 proves Property II in Theorem 4.5 and, in Theorem 4.4, gives a new short proof of Sibson’s theorem on angle vector optimality for the order-1 Delaunay triangulation [22]. It also discusses possible extensions to the class of maximal level-$2$ hypertriangulations and to levels beyond $2$. Section 5 proves Property III in Theorem 5.4, which extends it to order-$3$ for points in convex position in Theorem 5.5. Finally, Section 6 concludes the paper with discussions of open questions and conjectures related to the geometry and combinatorics of Delaunay and more general hypertriangulations.
@@ 2.1. Triangulations and Hypertriangulations
See Figure 1 for two level-$2$ hypertriangulations of five points. Observe that every triangulation of $A$ is a level-$1$ hypertriangulation of $A$, and vice versa, but for $k>1$, only a subset of the triangulations of $A^{(k)}$ are level-$k$ hypertriangulations of $A$. Note also that it is possible that a point can be written as the average of more than one subset of $k$ points in $A$: for example, the center of a square is the $2$-fold average of two pairs of diagonally opposite vertices. If a level-$k$ hypertriangulation uses such a point as a vertex, then it can use only one of the possible labels.
An alternative approach to these concepts is via induced subdivisions; see [24, Chapter 9] for details, including the definitions of induced subdivisions and tight subdivisions. According to this approach, a triangulation of $A=\{a_{1},a_{2},\ldots,a_{n}\}$ is a tight subdivision of ${{\rm conv\,}{A}}$ induced by the projection $\pi\colon\Delta_{n}\to{{\mathbb{R}}}^{2}$, in which $\Delta_{n}={{\rm conv\,}{\{e_{1},e_{2},\ldots,e_{n}\}}}\subseteq{{\mathbb{R}}}^{n}$ is the standard $(n-1)$-simplex, and $\pi(e_{i})=a_{i}$, for $i=1,2,\ldots,n$. To generalize, Olarte and Santos [17] use the level-$k$ hypersimplex, $\Delta_{n}^{(k)}$, which is the convex hull of the $k$-fold averages of the $e_{i}$ in ${{\mathbb{R}}}^{n}$, and define a level-$k$ hypertriangulation of $A$ as a tight subdivision of $A^{(k)}$ induced by the same projection $\pi$ restricted to $\Delta_{n}^{(k)}$. In this setting, the constraint to use only one label for each vertex is implicit.
@@ 2.2. The Aging Function
For example, the black triangles in the level-$2$ hypertriangulations in Figure 1 are obtained by aging the triangles in a complete triangulation on the left and a partial triangulation on the right. The aging function increases the level of the triangle by one, hence the name. Correspondingly, the inverse aging function maps a black triangle to a white triangle one level lower.
To extend this definition to hypertriangulations, we say a level-$k$ hypertriangulation, $P_{k}$, ages to a level-$(k+1)$ hypertriangulation, $P_{k+1}$, denoted $P_{k+1}=F(P_{k})$, if the aging function defines a bijection between the white triangles in $P_{k}$ and the black triangles in $P_{k+1}$. Note however that the aging of $P_{k}$ is not unique as it says nothing about the white triangles of $P_{k+1}$. This notion is useful to obtain structural results for the family of all level-$k$ hypertriangulations. For example, [8] has shown that every level-$2$ hypertriangulation is an aging of a level-$1$ hypertriangulation. For the special case in which the points are in convex position, [11] has extended this result to all levels, $k$. However, for points in possibly non-convex position, there are obstacles to applying the aging function. An example of a level-$2$ hypertriangulation, $P_{2}$, for which $F(P_{2})$ does not exist is given in [8, 17].
Refer to $c$$d$$e$$b$$c$$d$$e$$a$
Figure 1. Two level-$2$ hypertriangulations (drawn with solid segments) obtained by aging the triangulations (drawn with dotted segments) of the five white points, $a,b,c,d,e$, with four triangles on the left and two triangles on the right. Black triangles are shaded and white triangles are just white. The total number of triangles is the same in both hypertriangulations.
For later reference, we compile several results about the relation between level-$1$ and level-$2$ hypertriangulations obtained in [8]. Given a vertex, $x$, in a triangulation, $P$, we define the star of $x$ as the union of triangles that share $x$, denoted ${{\rm st}({P,x})}$, and shrinking the star by a factor two toward $x$, we get ${[{{{\rm st}({P,x})},x}]}=\frac{1}{2}({{\rm st}({P,x})}+x)$, which is the set of midpoints between $x$ and any point $y\in{{\rm st}({P,x})}$. Observe that the shrunken star is contained in ${{\rm conv\,}{A^{(2)}}}$ iff $x$ is an interior vertex of $P$. Indeed, $x$ necessarily belongs to the shrunken star, but if $x$ is a convex hull vertex, then $x$ lies outside ${{\rm conv\,}{A^{(2)}}}$. The first two statements of the following lemma are Lemmas 4.1 and 4.3 in [8], while the last statement is a reformulation of Lemma 4.2 in [8]. We refer to [8, Section 4] for proofs and additional details.
@@
Let $A\subseteq{{\mathbb{R}}}^{2}$ be finite and generic, and recall that every level-$1$ hypertriangulation is a triangulation.
@@ 2.3. Maximal and Complete Hypertriangulations
For example, the level-$2$ hypertriangulations in Figure 1 are both maximal, but only the one on the left is complete. The notion of maximality extends to level-$k$ hypertriangulations for any relevant $k$. The situation with completeness is more subtle as it relies on the existence of compatible triangulations to which we can apply the aging function, and there are counterexamples to the existence from level $2$ to level $3$; see Figure 8 in [8]—which is based on Example 5.1 in [17]—for a level-2 hypertriangulation with overlapping black triangles in level 3.
@@
As before, let $H\subseteq A$ be the vertices of the convex hull of $A$. To prove the first claim, let $n={{\#}{A}}$, $h={{\#}{H}}$, and consider a level-$2$ hypertriangulation, $P_{2}$, aged from a possibly partial triangulation, $P$, with $m+h\leq n$ vertices. Note that $P$ has $22m+h-$ 2 triangles, so $P_{2}$ has the same number of black triangles.
@@ 3.1. Higher Order Delaunay Triangulations
Figure 2. The (blue) order-$2$ Delaunay triangulation drawn on top of the (black) order-$2$ Voronoi tessellation of the set $A=\{a,b,\ldots,h\}$. Not all parts of the order-$2$ Voronoi tessellation are visible in the rectangular window.
@@ 3.2. Angles of Black and White Triangles
For $k=1$, there are no black triangles, so (bb) and (bw) are void. Delaunay [6] proved that the local angle property characterizes the (closest-point) Delaunay triangulation among all (complete) triangulations of a finite point set, and this was used by Lawson [14] to construct the triangulation by repeated edge flipping. Symmetrically, for $k={{\#}{A}}-1$, there are no white triangles, so (ww) and (bw) are void. Eppstein [9] proved the local angle property for the (farthest-point) Delaunay triangulation, and the convergence of the flip-algorithm implies that it is the only (not necessarily complete) triangulation of the points that has this property. The goal of this section is to extend these results to level-$k$ hypertriangulations.
@@ 3.3. All Delaunay Triangulations Have the Local Angle Property
We conjecture that among all level-$k$ hypertriangulations that satisfy the local angle property, the order-$k$ Delaunay triangulation is the only one that maximizes the number of triangles. For later reference, we refer to this as the Local Angle Conjecture for hypertriangulations. We discuss partial results for it in Section 5 and rigorously formulate it as Conjecture B in Section 6.
@@ 3.4. Constrained Delaunay Triangulations
Given a bounded polygonal region, $R$, it is always possible to find a triangulation, $P$, of its vertices (the endpoints of its edges) that contains all edges of the region. Hence, every triangle of $P$ lies either completely inside or completely outside the region. The restriction of $P$ to $R$ consists of the triangles inside $R$, and we call this restriction a triangulation of $R$. For some choices of $P$, the restriction to $R$ looks locally like the Delaunay triangulation, namely when every edge that passes through the interior of $R$ satisfies (ww). It is not difficult to see that such choices of triangulations exist and that their restriction to $R$ is generically unique: run Lawson’s algorithm on an initial triangulation of $R$, flipping an interior edge whenever the sum of the two opposite angles exceeds $\pi$. This is the constrained Delaunay triangulation of $R$, as introduced in 1989 by Paul Chew [4], but see also [3, 13]. A triangle $uvw$ belongs to this specific triangulation iff it is contained in $R$ and its circumcircle does not enclose any vertex that is visible from all points inside the triangle. Here, a point $x$ is visible from a point $y$, if the segment $xy$ lies inside $R$. We state a weaker necessary condition for later reference.
@@
Let $R$ be a bounded polygonal region in ${{\mathbb{R}}}^{2}$, assume its vertex set is generic, and let $u,v,w$ be vertices of $R$. If the triangle $uvw$ is contained in $R$, and its circumcircle does not enclose any vertex of $R$, then $uvw$ is a triangle in the constrained Delaunay triangulation of $R$. Similarly, if the segment $uv$ is contained in $R$ but is not an edge of $R$, and it has a circumcircle that does not enclose any vertex of $R$, then $uv$ is an edge of the constrained Delaunay triangulation of $R$.
We use constrained Delaunay triangulations to decompose white regions in aged hypertriangulations. To explain, let $P$ be a complete triangulation of a finite and generic set, $A\subseteq{{\mathbb{R}}}^{2}$, let $x\in A$ be a vertex of this triangulation, call ${{\rm wh}({P,x})}={{\rm st}({P,x})}\cap{{\rm conv\,}{(A\setminus\{x\})}}$ the white region of $x$ in $P$, and let $P(x)$ be a triangulation of ${{\rm wh}({P,x})}$. Note that ${{\rm wh}({P,x})}={{\rm st}({P,x})}$ if $x$ is an interior vertex, and ${{\rm wh}({P,x})}\subsetneq{{\rm st}({P,x})}$ if $x$ is a convex hull vertex. In the special case in which $P$ is the order-$1$ Delaunay triangulation and $P(x)$ is the constrained Delaunay triangulation of ${{\rm wh}({P,x})}$, for each $x\in A$, these sets contain all white triangles in the order-$2$ Delaunay triangulation, albeit the latter are only half the size. This can be seen from the following observation. If we consider the (order-1) Delaunay triangulation of $A\setminus\{x\}$, then it contains old triangles from $P$ not incident to $x$ and new triangles that retriangulate ${{\rm st}({P,x})}\cap{{\rm conv\,}{(A\setminus\{x\})}}$. The circumcircle of each new triangle encloses $x$. Thus, the new triangles are the white triangles from Definition 3.1 and also the triangles from the constrained Delaunay triangulation of ${{\rm st}({P,x})}\cap{{\rm conv\,}{(A\setminus\{x\})}}$.
More generally, we use the constrained Delaunay triangulations of the white regions to disambiguate the aging function. This is done extensively in the proofs of our main results in Section 4, where they are useful because constrained Delone triangulations optimize angle vectors, and in Section 5, where they are useful because they satisfy the local angle property.
@@ 4.1. Triangulations and Angle Vectors
According to Lemma 2.6, all maximal level-$2$ hypertriangulations have the same number of triangles. We can therefore compare their angle vectors lexicographically as described.
@@ 4.2. Elementary Lemmas
Assume without loss of generality that $L$ is horizontal, and let $A^{\prime}\subseteq A$ contain all points strictly above $L$. The boundary of ${{\rm conv\,}{A}}$ is a closed convex curve, $\gamma$, and we write $\gamma^{\prime}\subseteq\gamma$ for the vertices and edges strictly above $L$. Every point $a\in A^{\prime}$ is either a vertex of $\gamma^{\prime}$, or there is an edge $ab$ in $P$, with $b$ above $L$ and farther from $L$ than $a$. Hence, $ab\in Q$. We can therefore trace a path from $a$ that eventually reaches a vertex of $\gamma^{\prime}$ in $Q$, which implies that the part of $Q$ strictly above $L$ is either empty or connected. Symmetrically, the part of $Q$ strictly below $L$ is either empty or connected, which implies the claim. ∎
@@
    If $v=b$ is the only shared vertex between $abc$ and $uvw$, then $uw$ cannot cross both $ab$ and $bc$.